OpenStudy (anonymous):

. Dale and Della are making Kool-Aid. Dale mixes a 20-ounce drink composed of 25% sugar while Della mixes an 18-ounce drink with 30% sugar. If the two drinks are combined, what percent (rounded to the nearest percent) of the new mixture is composed of sugar? please show me your work.

OpenStudy (matt101):

First find the amount of sugar you're dealing with (if you have experience in physical chem, you'll know this is similar to combining mixtures of different concentrations). 0.25*20=5 oz of sugar in the first drink 0.30*18=5.4 oz of sugar in the second drink 5+5.4 = 10.4 oz of sugar total When the drinks are mixed, you're dealing with 20+18=38 oz of drink. 10.4 oz are sugar, so as a percent, 10.4/38 * 100 = 27.4%

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

thanks for explaining